Fran receives $153 per year in simple interest from investing $1900. Part is invested at 7%, 8%, and 9%. There is a $100 more invested at 9% than 8%. find the amount invested at each rate
Answer by TimothyLamb(4379)

x = amount invested at 7%
y = amount invested at 8%
z = amount invested at 9%
---
x + y + z = 1900
0.07x + 0.08y + 0.09z = 153
z = y + 100
---
put the system of linear equations into standard form
---
x + y + z = 1900
0.07x + 0.08y + 0.09z = 153
y - z = -100
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = amount invested at 7% = 600
y = amount invested at 8% = 600
z = amount invested at 9% = 700
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---